2/5 of the sum of two numbers is is 18. one of the numbers is 30, what is the other number?
Answer by rm29924(97)


Let x=the other number
2/5(x+30)=18
2/5x+2/5(30)=18
2/5x+12=18
2/5x=6
multiply both sides by 5 to get rid of the fraction
2x=30
X=15